# A Generalized Adaptive Joint Learning Framework for High-Dimensional Time-Varying Models

### v1

@@ Abstract
In modern biomedical and econometric studies, longitudinal processes are often characterized by complex time-varying associations and abrupt regime shifts that are shared across correlated outcomes. Standard functional data analysis (FDA) methods, which prioritize smoothness, often fail to capture these dynamic structural features, particularly in high-dimensional settings. This article introduces Adaptive Joint Learning (AJL), a regularization framework designed to simultaneously perform functional variable selection and structural changepoint detection in multivariate time-varying coefficient models.We propose a convex optimization procedure that synergizes adaptive group-wise penalization with fused regularization, effectively borrowing strength across multiple outcomes to enhance estimation efficiency. We provide a rigorous theoretical analysis of the estimator in the ultra-high-dimensional regime ($p\gg n$), establishing non-asymptotic error bounds and proving that AJL achieves the oracle property—performing as well as if the true active set and changepoint locations were known a priori. A key theoretical contribution is the explicit handling of approximation bias via undersmoothing conditions to ensure valid asymptotic inference. The proposed method is validated through comprehensive simulations and an application to Primary Biliary Cirrhosis (PBC) data. The analysis reveals previously characterizing synchronized phase transitions in disease progression and identifying a parsimonious set of time-varying prognostic markers.
@@ 1 Introduction
However, several critical gaps remain in the literature. Standard $\ell_{1}$ and group-$\ell_{1}$ penalties (like the Group Lasso, Yuan and Lin (2006)) are known to produce biased estimates for large coefficients, thus failing to possess the oracle property (Fan and Li, 2001). This property, which ensures an estimator is as asymptotically efficient as if the true sparse model were known in advance, is crucial for accurate inference. This limitation can be overcome by using adaptive weights, as in the Adaptive Lasso (Zou, 2006) and its extensions (Wang and Leng, 2008), or by using non-convex penalties like SCAD (Fan and Li, 2001) or MCP (Zhang, 2010). Furthermore, standard functional models enforce smoothness (e.g., via $\int[\beta^{\prime\prime}(t)]^{2}dt$). In many longitudinal applications, however, we are interested in abrupt changepoints, for instance, a sudden change in disease trajectory or symptom onset. The fused lasso (Tibshirani et al., 2005) and its variants (Rinaldo, 2009; Viallon et al., 2016) are designed for this, but are rarely integrated with high-dimensional functional variable selection. Finally, in many studies (e.g., clinical trials, neuroimaging), multiple correlated outcomes ($K>1$) are collected. Analyzing each outcome separately is statistically inefficient. Joint modeling, or multi-task learning, can borrow strength across outcomes to improve estimation and selection accuracy (Obozinski et al., 2010; Caruana, 1997; Argyriou et al., 2008).
@@
We develop a scalable multi-stage BCD-ADMM algorithm, provide a rigorous theoretical analysis with detailed proofs, and validate our method via extensive, redesigned simulation studies. Finally, we demonstrate the applicability of AJL by analyzing the classic Mayo Clinic Primary Biliary Cirrhosis (PBC) dataset. We jointly model the trajectories of three key liver function markers, identifying shared structural breakpoints in disease progression and revealing how the prognostic effects of baseline risk factors evolve dynamically over a decade of follow-up.
The remainder of this paper is organized as follows. Section 2 introduces the functional TVCM, the basis expansion, and the motivations for our penalized objective. Section 3 details the optimization procedure, including convergence and complexity analysis. Section 4 presents the core theoretical results, including the high-dimensional assumptions, non-asymptotic error bounds, and asymptotic oracle properties. Section 5 validates our method through extensive numerical studies. Section 6 provides a detailed application to PBC data analysis. Section 7 concludes. The Appendix contains proofs and discusses potential extensions.
@@ 2.1 The Time-Varying Coefficient Model (TVCM)
While baseline covariates are prevalent in clinical studies, modern longitudinal designs (e.g., wearable sensors, intensive monitoring) often yield densely sampled predictor processes. To accommodate such complexity, we propose a general Concurrent Varying-Coefficient Model (TVCM) framework:
@@
where $\bm{\alpha}(t)=(\alpha_{1}(t),\dots,\alpha_{K}(t))^{\top}$ represents the $K$ functional intercepts, and $\bm{\beta}_{k}(t)=(\beta_{1k}(t),\dots,\beta_{pk}(t))^{\top}$ denotes the vector of $p$ time-varying functional coefficients for the $k$-th outcome. The errors $\epsilon_{ilk}$ are assumed to be independent with mean zero and finite variance.
@@ 2.5 Theoretical Motivation for Adaptivity
2.6 The Generalized Adaptive Joint Learning (AJL) Objective
@@
$Q(\theta)=$ $\frac{1}{2N}\left\|Y-Z_{\Phi}A-\sum_{j=1}^{p}X_{\Phi,j}B_{j}\right\|_{F}^{2}+\lambda{g}\sum_{j=1}^{p}\hat{w}_{g,j}\|B_{j}\|_{F}\quad$ $(\text{Selection})$ (10)
$+\lambda{f}^{\alpha}\sum_{k=1}^{K}\sum_{m=1}^{M-1}\hat{w}_{f,km}^{\alpha}|a_{k,m+1}-a_{k,m}|\quad$ $(\text{Intercept Shifts})$
@@
    Adaptive Slope Fusion: The term with $\lambda_{f}^{\beta}$ enforces piecewise-constant dynamics in the covariate effects $\beta_{jk}(t)$. This allows the model to identify critical time points where the relationship between a predictor and the outcome undergoes a structural shift (e.g., a ”regime change” in treatment efficacy).
@@
Remark 2 (Hierarchical Regularization Strategy and Computational Strategy).
While the generalized objective (10) offers a holistic solution, simultaneously optimizing for selection and slope-segmentation in the ultra-high-dimensional regime ($p\gg n$) presents computational challenges and potential identification issues due to the interaction between the group-sparsity and fusion penalties.
To address this, we adopt a Hierarchical Regularization Strategy motivated by the principle of Sure Screening (Fan and Lv, 2008).
Screening Phase (Focus of this paper): We set $\lambda_{f}^{\beta}=0$ (or sufficiently small) to prioritize the consistent recovery of the active support set $\mathcal{S}$ via the Adaptive Group Lasso, while simultaneously detecting major baseline shifts via $\lambda_{f}^{\alpha}$. This ensures that the dimensionality is effectively reduced from $p$ to $s=|\mathcal{S}|\ll n$.
Refinement Phase: Conditional on the selected support $\mathcal{S}$, one can subsequently activate the slope fusion penalty ($\lambda_{f}^{\beta}>0$) to refine the structural dynamics of the non-zero effects.
Our theoretical analysis in Section 4 demonstrates that the Screening Phase achieves the Oracle Property for variable selection. This implies that the separation of stages does not compromise the asymptotic validity of the estimator, but rather provides a numerically robust path to the global oracle solution. Consequently, our simulation studies and algorithm primarily focus on verifying the crucial Screening Phase.
While the objective in (10) is convex, solving it for ultra-high-dimensional $p$ with simultaneous Group and Fused penalties on $B_{j}$ can be computationally intensive. In practice, we often adopt a hierarchical strategy: we first set $\lambda_{f}^{\beta}=0$ to identify the active support set $\mathcal{S}$ via the Adaptive Group Lasso. Subsequently, we solve the full objective restricted to $j\in\mathcal{S}$ to refine the estimation of temporal dynamics. This two-stage procedure consistently recovers the oracle structure under the assumption that signal strength satisfies the ”beta-min” condition, effectively reducing the computational complexity from $O(p)$ to $O(|\mathcal{S}|)$ for the fusion step.
@@ 4 Theoretical Analysis
Unlike prior works that treat structural discovery and variable selection in isolation, our theoretical analysis establishes the oracle properties for the comprehensive Generalized AJL objective (10), fully incorporating the slope-fusion regularization component. Leveraging the block-decomposability of B-spline coefficients and the structural isomorphism between intercept and slope difference penalties, we derive the consistency and asymptotic normality for the full estimator. Furthermore, we explicitly provide the theoretical justification for the Hierarchical Regularization Strategy (Remark 2). By establishing the Sure Screening Property of the initial stage, we guarantee that the computationally efficient screening phase consistently recovers the true active set, thereby validating the asymptotic optimality of the sequential refinement procedure.
@@ 4.2 Non-Asymptotic Analysis
Let $\bm{\theta}=\text{vec}(\bm{A},\bm{B}_{1},\dots,\bm{B}_{p})$ be the vector of all coefficients. Following the modern framework for high-dimensional M-estimators (Bühlmann and van de Geer, 2011; Negahban et al., 2012), we first establish non-asymptotic error bounds.
@@
Proposition 2 (Validity of Hierarchical Screening).
@@
Remark 4.
Proposition 2 provides the theoretical justification for our Hierarchical Regularization Strategy. It guarantees that, with probability approaching one, the initial screening step retains all truly relevant predictors. Consequently, the subsequent refinement step (where $\lambda_{f}^{\beta}>0$ is activated) operates within a subspace that contains the true model, ensuring that the final structural discovery is asymptotically unbiased.
The proofs of Lemma 1, Theorems 1 and 2 are non-trivial and build on extensions of the general frameworks in Negahban et al. (2012) and Bühlmann and van de Geer (2011) to our adaptive fused-group structure; detailed arguments are provided in Appendix A.
4.3 Asymptotic Oracle Property
Theorem 3 (Asymptotic Selection Consistency).
Let $(\widehat{\bm{A}},\{\widehat{\bm{B}}_{j}\})$ be the solution to the AJL problem (10). Under Assumptions 1-6, the estimator is selection-consistent:
  1. 1.
    $\text{Pr}(\{j:\|\widehat{\bm{B}}_{j}\|_{F}>0\}=\mathcal{S})\to 1$ as $N\to\infty$.
  2. 2.
    $\text{Pr}(\left\{m:\hat{a}_{k,m+1}\neq\hat{a}_{k,m}\right\}=\mathcal{C}_{k},\forall k)\to 1$ as $N\to\infty$.
Theorem 4 (Asymptotic Normality and Valid Inference).
Suppose Assumptions A1-A6 hold, and additionally, the Undersmoothing Condition (Assumption 7) is satisfied (i.e., $NM_{N}^{-2d}\to 0$ where $d$ is the spline smoothness order). Let $\mathcal{A}=\mathcal{S}\cup\mathcal{C}^{\alpha}\cup\mathcal{C}^{\beta}$ denote the Global Oracle Set, comprising the true active predictors, baseline changepoints, and slope dynamics. Let $\hat{\bm{\theta}}_{\mathcal{A}}$ be the vector of estimated coefficients restricted to this set. Then, the estimator satisfies:
  1. 1.
    Oracle Distribution of Coefficients: The penalized estimator behaves asymptotically as the unpenalized Oracle estimator on the true support:
    where $\bm{\theta}_{\mathcal{A}}^{*}$ is the projection of the true functions onto the B-spline space, and $\bm{\Sigma}_{\mathcal{A}}=\sigma_{\epsilon}^{2}(\mathbb{E}[\mathbf{Z}_{\mathcal{A}}^{\top}\mathbf{Z}_{\mathcal{A}}])^{-1}$ is the asymptotic covariance matrix of the Oracle sub-model.
  2. 2.
    Pointwise Functional Inference: For any active predictor $j\in\mathcal{S}$ and any time point $t\in\mathcal{T}$ that is not a changepoint, the estimated time-varying coefficient $\hat{\beta}_{j}(t)$ follows:
    where $\beta_{j}^{*}(t)$ is the true functional curve (not just the spline approximation). The standard error $\hat{\sigma}_{j}(t)$ is consistently estimated via the plug-in estimator:
    $\hat{\sigma}_{j}^{2}(t)=\bm{\phi}(t)^{\top}[\widehat{\text{Cov}}(\hat{\bm{\theta}}_{\mathcal{A}})]_{jj}\bm{\phi}(t).$
The significance of (18) lies in the validity of post-selection inference. The Undersmoothing Condition ensures that the squared approximation bias is asymptotically negligible compared to the variance ($bias^{2}\ll var$). Consequently, we can construct valid $(1-\alpha)$ confidence intervals for the dynamic effects as $\hat{\beta}_{j}(t)\pm z_{\alpha/2}\hat{\sigma}_{j}(t)$, treating the selected structure as fixed.
@@ 4.3 Asymptotic Oracle Property
While recent high-dimensional inference literature has favored de-biasing techniques, e.g., Javanmard and Montanari (2014); van de Geer et al. (2014), to circumvent undersmoothing assumptions, extending these methods to the AJL framework presents unique and non-trivial challenges. First, from a computational perspective, the intrinsic high collinearity of B-spline basis functions (Assumption 5) renders the empirical Hessian matrix ill-conditioned. Constructing a stable approximate inverse (e.g., via nodewise regression) for the block-structured design matrix $Z_{\Phi}$ is numerically precarious compared to standard i.i.d. Gaussian designs.Second, and more fundamentally, theoretical de-biasing requires constructing a “decorrelated score function” that projects the noise away from the tangent space of the penalty. In our multi-task setting, the penalty $P_{ada}$ induces a complex geometry due to the structural non-orthogonality between the group-sparsity manifold and the piecewise-constant manifold. Developing a rigorous projection argument that simultaneously decorrelates against both functional selection bias and changepoint localization bias remains an open mathematical problem.In contrast, our approach leverages the Oracle Property (Theorem 3) to asymptotically eliminate the regularization bias on the active set. Under this regime, the remaining error is dominated by approximation bias. Consequently, the undersmoothing condition (Assumption 7) serves as a theoretically sufficient and standard functional data analytic tool to ensure valid inference without the instability of inverting high-dimensional spline matrices.
@@
A proof for Theorems 3 and 4 is provided in Appendix A. ∎

### v2

@@ Abstract
In modern biomedical and econometric studies, longitudinal processes are often characterized by complex time-varying associations and abrupt regime shifts that are shared across correlated outcomes. Standard functional data analysis (FDA) methods, which prioritize smoothness, often fail to capture these dynamic structural features, particularly in high-dimensional settings. This article introduces Adaptive Joint Learning (AJL), a hierarchical regularization framework designed to integrate functional variable selection with structural changepoint detection in multivariate time-varying coefficient models. Unlike standard simultaneous estimation approaches, we propose a theoretically grounded two-stage screening-and-refinement procedure. This framework first synergizes adaptive group-wise penalization with sure screening principles to robustly identify active predictors, followed by a refined fused regularization step that effectively borrows strength across multiple outcomes to detect local regime shifts. We provide a rigorous theoretical analysis of the estimator in the ultra-high-dimensional regime ($p\gg n$). Crucially, we establish the sure screening consistency of the first stage, which serves as the foundation for proving that the refined estimator achieves the oracle property—performing as well as if the true active set and changepoint locations were known a priori. A key theoretical contribution is the explicit handling of approximation bias via undersmoothing conditions to ensure valid asymptotic inference. The proposed method is validated through comprehensive simulations and an application to Sleep-EDF data, revealing novel dynamic patterns in physiological states.
@@ 1 Introduction
However, several critical gaps remain in the literature. Standard $\ell_{1}$ and group-$\ell_{1}$ penalties (like the Group Lasso, Yuan and Lin (2006)) are known to produce biased estimates for large coefficients, thus failing to possess the oracle property (Fan and Li, 2001). This property, which ensures an estimator is as asymptotically efficient as if the true sparse model were known in advance, is crucial for accurate inference. This limitation can be overcome by using adaptive weights, as in the Adaptive Lasso (Zou, 2006) and its extensions (Wang and Leng, 2008), or by using non-convex penalties like SCAD (Fan and Li, 2001) or MCP (Zhang, 2010). Furthermore, standard functional models enforce smoothness (e.g., via $\int[\beta^{\prime\prime}(t)]^{2}dt$). In many longitudinal applications, however, we are interested in abrupt changepoints, for instance, a sudden change in disease trajectory or symptom onset. While recent advances have developed efficient algorithms for multiple changepoint detection in high-dimensional generalized linear models (Wang et al., 2023) and quantile regression settings (Wang et al., 2025), these methods are primarily tailored for discrete time series or specific distributional parameters. They are not directly applicable to the functional data setting, where the goal is to detect structural breaks within continuous time-varying coefficient curves $\beta_{j}(t)$ while simultaneously performing variable selection. The fused lasso (Tibshirani et al., 2005) and its variants (Rinaldo, 2009; Viallon et al., 2016) are designed for this, but are rarely integrated with high-dimensional functional variable selection. Finally, in many studies (e.g., clinical trials, neuroimaging), multiple correlated outcomes ($K>1$) are collected. Analyzing each outcome separately is statistically inefficient. Joint modeling, or multi-task learning, can borrow strength across outcomes to improve estimation and selection accuracy (Obozinski et al., 2010; Caruana, 1997; Argyriou et al., 2008).
@@
We develop a scalable multi-stage BCD-ADMM algorithm, provide a rigorous theoretical analysis with detailed proofs, and validate our method via extensive, redesigned simulation studies. Finally, we demonstrate the applicability of AJL by analyzing the Sleep-EDF database. We jointly model the spectral power trajectories of multiple EEG frequency bands, identifying shared structural breakpoints that align with NREM-REM sleep cycles. Crucially, by leveraging the derived undersmoothing properties, we construct pointwise confidence bands to validate the time-varying effects of age and gender while rigorously filtering out high-dimensional noise variables.
The remainder of this paper is organized as follows. Section 2 introduces the functional TVCM, the basis expansion, and the motivations for our penalized objective. Section 3 details the optimization procedure, including convergence and complexity analysis. Section 4 presents the core theoretical results, including the high-dimensional assumptions, non-asymptotic error bounds, and asymptotic oracle properties. Section 5 validates our method through extensive numerical studies. Section 6 provides a application to EEG data analysis. Section 7 concludes. The Appendix contains proofs and discusses potential extensions.
@@ 2.1 The Time-Varying Coefficient Model (TVCM)
While baseline covariates are prevalent in clinical studies, modern longitudinal designs (e.g., wearable sensors, intensive monitoring) often yield densely sampled predictor processes. To accommodate such complexity, we propose a general concurrent Time-Varying Coefficient Model (TVCM) framework:
@@
where $\bm{\alpha}(t)=(\alpha_{1}(t),\dots,\alpha_{K}(t))^{\top}$ and $\bm{\beta}_{k}(t)=(\beta_{1k}(t),\dots,\beta_{pk}(t))^{\top}$ represents the $K$ functional intercepts and the vector of $p$ time-varying functional coefficients for the $k$-th outcome, respectively. The errors $\epsilon_{ilk}$ are assumed to be independent with mean zero and finite variance.
@@ 2.5 Theoretical Motivation for Adaptivity
2.6 The Two-Stage Generalized Adaptive Joint Learning (AJL) Framework
@@
$Q(\theta)$ $=\frac{1}{2N}\|Y-Z_{\Phi}A-\sum_{j=1}^{p}X_{\Phi,j}B_{j}\|_{F}^{2}+\lambda_{g}\sum_{j=1}^{p}\hat{w}_{g,j}\|B_{j}\|_{F}\quad$ $(\text{Selection})$ (10)
$+\lambda_{f}^{\alpha}\sum_{k=1}^{K}\sum_{m=1}^{M-1}\hat{w}_{f,km}^{\alpha}|a_{k,m+1}-a_{k,m}|\quad$ $(\text{Intercept Shifts})$
@@
    Adaptive Slope Fusion: The term with $\lambda_{f}^{\beta}$ enforces piecewise-constant dynamics in the covariate effects $\beta_{jk}(t)$. This allows the model to identify critical time points where the relationship between a predictor and the outcome undergoes a structural shift (e.g., a “regime change” in treatment efficacy).
@@
However, directly minimizing (10) in the ultra-high-dimensional regime ($p\gg n$) presents two fundamental challenges. First, from a computational perspective, simultaneously enforcing group sparsity and slope fusion for all $p$ predictors is computationally prohibitive. Second, theoretically, the structural interference between the covariate selection space (governed by the group lasso) and the temporal segmentation space (governed by the fused lasso) can lead to identification issues when the signal-to-noise ratio is low.
To address these challenges, we propose a constructive Two-Stage Screening-and-Refinement Framework. Instead of a single-step optimization, we adopt a hierarchical strategy motivated by the principle of Sure Screening (Fan and Lv, 2008). This framework decomposes the problem into a screening phase that prioritizes support recovery, followed by a refinement phase that focuses on structural discovery.
In the first stage, we define the Screening Estimator $\hat{\theta}^{scr}$ by simplifying the global objective to prioritize support recovery. Specifically, we set $\lambda_{f}^{\beta}=0$ (enforcing no slope segmentation initially) and solve:
[…]
Let $\hat{\mathcal{S}}_{scr}=\{j:\|\hat{B}_{j}^{scr}\|_{F}>0\}$ denote the selected active set. This stage serves to reduce the dimensionality from $p$ to $|\hat{\mathcal{S}}_{scr}|\ll n$.
In the second stage, we obtain the Refined Estimator $\hat{\theta}^{ref}$ by optimizing the full objective (including the slope fusion penalty $\lambda_{f}^{\beta}$) restricted to the selected subspace $\hat{\mathcal{S}}_{scr}$:
$\hat{\theta}^{ref}=\arg\min_{\theta:B_{j}=0,\forall j\notin\hat{\mathcal{S}}_{scr}}\left(L_{N}(\theta)+P_{ada}(\theta)\right)$
This refinement step allows for the precise localization of dynamic changepoints within the identified relevant predictors.
@@ 4 Theoretical Analysis
Unlike prior works that treat structural discovery and variable selection in isolation, our theoretical analysis establishes the oracle properties for the comprehensive Generalized AJL objective (10), fully incorporating the slope-fusion regularization component. Leveraging the block-decomposability of B-spline coefficients and the structural isomorphism between intercept and slope difference penalties, we derive the consistency and asymptotic normality for the full estimator. Furthermore, we explicitly provide the theoretical justification for the Hierarchical Regularization Strategy. By establishing the Sure Screening Property of the initial stage, we guarantee that the computationally efficient screening phase consistently recovers the true active set, thereby validating the asymptotic optimality of the sequential refinement procedure.
@@ 4.2 Non-Asymptotic Analysis
Since our computational strategy relies on a hierarchical screening procedure, our first theoretical priority is to establish the validity of this reduction. The following theorem guarantees that the initial screening stage retains all relevant functional predictors with probability approaching one.
Theorem 1 (Validity of Hierarchical Screening).
[…]
Remark 3.
Theorem 1 provides the theoretical justification for our Hierarchical Regularization Strategy. It guarantees that, with probability approaching one, the initial screening step retains all truly relevant predictors. Consequently, the subsequent refinement step (where $\lambda_{f}^{\beta}>0$ is activated) operates within a subspace that contains the true model, ensuring that the final structural discovery is asymptotically unbiased.
Conditioned on the sure screening property established in Theorem 1, we now investigate the asymptotic properties of the refined estimator. The following results demonstrate that, by optimizing the generalized objective over the selected subspace, the AJL estimator achieves optimal estimation rates and consistent structural recovery.
Let $\bm{\theta}=\text{vec}(\bm{A},\bm{B}_{1},\dots,\bm{B}_{p})$ be the vector of all coefficients. Following the modern framework for high-dimensional M-estimators (Bühlmann and van de Geer, 2011; Negahban et al., 2012), we establish non-asymptotic error bounds.
@@
The proofs of Lemma 1, Theorems 2 and 3 are non-trivial and build on extensions of the general frameworks in Negahban et al. (2012) and Bühlmann and van de Geer (2011) to our adaptive fused-group structure; detailed arguments are provided in Appendix A.
4.3 Asymptotic Oracle Property
Beyond point estimation and selection, valid statistical inference is crucial for scientific discovery. Building on the selection consistency (Theorem 3), we now establish the asymptotic normality of the functional estimator, justifying the construction of pointwise confidence bands.
Theorem 4 (Structural Selection Consistency).
Suppose Assumptions 16 hold. Given the sure screening property established in Theorem 1 (i.e., the true active set is retained in the first stage), the Refined Estimator $\hat{\bm{\theta}}^{ref}$ achieves consistent structural recovery. Specifically, as $N\to\infty$:
  1. 1.
    Variable Selection Consistency: The estimator correctly identifies the true functional support:
    $\text{Pr}\left(\{j:\|\hat{B}^{ref}_{j}\|_{F}>0\}=\mathcal{S}^{*}\right)\to 1.$ (17)
  2. 2.
    Changepoint Detection Consistency: The estimator correctly locates the structural breaks in both baseline and time-varying effects:
    $\text{Pr}\left(\hat{\mathcal{C}}^{\alpha}=\mathcal{C}^{\alpha*}\right)\to 1,$ (18)
    $\text{Pr}\left(\hat{\mathcal{C}}^{\beta}_{j}=\mathcal{C}^{\beta*}_{j},\forall j\in\mathcal{S}^{*}\right)\to 1,$ (19)
    where $\hat{\mathcal{C}}^{\alpha}=\{m:\hat{a}_{k,m+1}^{ref}\neq\hat{a}_{k,m}^{ref}\}$ and $\hat{\mathcal{C}}^{\beta}_{j}=\{m:\hat{b}_{jk,m+1}^{ref}\neq\hat{b}_{jk,m}^{ref}\}$ denote the sets of estimated jump locations.
Theorem 5 (Asymptotic Normality and Valid Inference).
Suppose Assumptions 1-6 hold. Additionally, assume the undersmoothing condition (Assumption 7) is satisfied (i.e., $NM_{N}^{-2d}\to 0$, where $d$ is the spline smoothness order). Given the sure screening property established in Theorem 1, the refined estimator $\hat{\theta}^{ref}$ effectively operates on the true support with probability approaching one.
Let $\mathcal{I}_{\mathcal{A}}\subset\{1,\dots,pM_{N}+KM_{N}\}$ denote the Oracle Index Set, representing the indices of non-zero B-spline coefficients corresponding to the true active predictors and their structural segments. Let $\hat{\bm{\theta}}_{\mathcal{A}}$ denote the sub-vector of estimated coefficients restricted to this set.
  1. 1.
    Oracle Distribution of Coefficients: The penalized estimator on the active set is asymptotically equivalent to the unpenalized Oracle estimator, satisfying:
    where $\bm{\theta}_{\mathcal{A}}^{*}$ is the projection of the true functions onto the B-spline space, and $\bm{\Sigma}_{\mathcal{A}}$ is the asymptotic covariance matrix of the Oracle sub-model.
  2. 2.
    Pointwise Functional Inference: For any active predictor $j\in\mathcal{S}^{*}$ and any time point $t\in\mathcal{T}$ (excluding changepoints), the estimated time-varying coefficient $\hat{\beta}_{j}(t)$ follows:
    where $\beta_{j}^{*}(t)$ is the true functional curve. The standard error $\hat{\sigma}_{j}(t)$ is consistently estimated via:
    $\hat{\sigma}_{j}^{2}(t)=\bm{\phi}(t)^{\top}\widehat{\bm{\Omega}}_{jj}\bm{\phi}(t),$ (22)
    where $\widehat{\bm{\Omega}}_{jj}$ denotes the $M_{N}\times M_{N}$ diagonal block of the estimated covariance matrix $\widehat{\text{Cov}}(\hat{\bm{\theta}}_{\mathcal{A}})$ corresponding to the $j$-th predictor.
Remark 4 (Validity of Post-Selection Inference).
@@ 4.3 Asymptotic Oracle Property
While recent high-dimensional inference literature has favored de-biasing techniques, e.g., Javanmard and Montanari (2014); van de Geer et al. (2014), to circumvent undersmoothing assumptions, extending these methods to the AJL framework presents unique and non-trivial challenges. First, from a computational perspective, the intrinsic high collinearity of B-spline basis functions (Assumption 5) renders the empirical Hessian matrix ill-conditioned. Constructing a stable approximate inverse (e.g., via nodewise regression) for the block-structured design matrix $Z_{\Phi}$ is numerically precarious compared to standard i.i.d. Gaussian designs.Second, and more fundamentally, theoretical de-biasing requires constructing a “decorrelated score function” that projects the noise away from the tangent space of the penalty. In our multi-task setting, the penalty $P_{ada}$ induces a complex geometry due to the structural non-orthogonality between the group-sparsity manifold and the piecewise-constant manifold. Developing a rigorous projection argument that simultaneously decorrelates against both functional selection bias and changepoint localization bias remains an open mathematical problem. In contrast, we adopt the undersmoothing strategy, which is the standard paradigm for valid inference in semi-parametric sieve estimation (Chen, 2007). As rigorously established by Shen et al. (1998) for regression splines and generalized by Newey (1997) for series estimators, valid confidence bands can be constructed by allowing the number of knots to grow at a rate where the squared approximation bias becomes asymptotically negligible relative to the variance ($bias^{2}\ll var$). In our high-dimensional context, conditioned on the oracle selection consistency (Theorem 4), our estimator behaves asymptotically as a series estimator on the active set. Consequently, the undersmoothing condition, Assumption 7 serves as a theoretically sufficient condition to center the asymptotic distribution at the true function $\beta^{*}(t)$ without the numerical instability of inverting high-dimensional spline matrices.
@@
A proof for Theorems 4 and 5 is provided in Appendix A. ∎